OpenStudy (anonymous):

please show all work -4y+10=-2(3y+1)

OpenStudy (dumbcow):

y=-6

OpenStudy (anonymous):

\[y=-6\] \[(-4y+10)/-2 = (-2(3y+1))/-2\] 2y-5=3y+1 2y-5+5=3y+1+5 2y=3y+6 2y-3y=6 -y=6 y=-6